OpenStudy (anonymous):

A 3 digit natural number is selected at random .Find the probability that it is divisible by 2 but not by 3?

OpenStudy (anonymous):

y u delete the answer

OpenStudy (anonymous):

I think I missed something

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so by solving it , we can get asnwer

OpenStudy (anonymous):

2/3 of (9*10*5)/(9*10*10)

OpenStudy (anonymous):

So, it is 1/3

OpenStudy (anonymous):

final answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

see we have 3 digit nos from 100 to 999 and we have to find prob , no should div by 2 not by 3

OpenStudy (anonymous):

i think something is missing

OpenStudy (anonymous):

number should be divisible by 2 and not by 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ExperimentX would u plse guide me

OpenStudy (anonymous):

So, there are 9*10*5, 3 digit numbers that are divisible by 2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@Shubhamsrg

OpenStudy (shubhamsrg):

nos divisible by 2= 100,102...998 ->450 nos. nos divisible by 6(both 2 and 3) = 102,108..996 ->150 nos. so total nos we have to take care of = 450 - 150 = 300 total nos. from 100 to 999 = 900 so our required probability = 300/900 = 1/3 as @sauravshakya rightly said.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thank s